What is the number of moles in 500 L of He gas at STP?
1
Expert's answer
2012-12-11T08:44:51-0500
PV=nRT

n=PV/RT ,but under STP& P/RT=1/22.4

So using this,or empirical equation below:

n & 500L
He -------> He
1 & 22.4

n=500/22.4=22.321 moles